### v1

@@ Abstract
Figure 1: Comparison of different paradigms. (a) Existing methods fully fine-tune the feature extractor and combine it with temporal alignment module. (b) We froze the image encoder and apply Task-Specific adaptation to extract the most discriminative information. (c) We use semactic order adapter to model sequential information within sub-actions and further calculate the score by stage with aligned visual features.
@@ 1 Introduction
With the advent of large-scale pre-trained models [31, 19], researchers are now able to exploit the knowledge these models have acquired during pre-training and apply it to downstream tasks. Accordingly, there is an increasing focus on fine-tuning large pre-trained models for specific tasks [10, 63, 54, 49]. In video understanding, existing works typically fine-tune pre-trained models on well-labeled video datasets [5, 11], where the training and testing categories are identical. However, in realistic scenarios, collecting sufficient and well-labeled video samples is inherently challenging. As a result, few-shot learning has garnered increasing attention [29, 6, 37, 55, 2]. A common practice in few-shot action recognition is to adopt a metric-based meta-learning strategy, where classification probabilities are computed based on the distance between query video features and support prototypes. Prior works [26, 4, 30, 43, 14, 20] achieve this by fine-tuning large image models, such as CLIP [31], on video datasets alongside independent temporal modules, as shown in Fig. 1 (a). Beyond visual-centric approaches, recent studies have explored multimodal frameworks for few-shot learning by integrating semantic information. For instance, works like [45, 17, 57] employ LLM or VLM to expand label descriptions, demonstrating broad utility.
@@
Although previous studies have made progress, several limitations remain. First, full fine-tuning erodes the model’s inherent prior knowledge, thereby degrading its ability to generalize to unseen tasks. Second, the importance of incorporating task-specific adaptation into parameter-efficient fine-tuning (PEFT) is often overlooked, which leads to the neglect of the most discriminative information in each meta-task. Third, enhanced label information is typically treated as a static description, ignoring dynamic and ordered aspects that are essential for capturing semantic cues in fine-grained sequential modeling. Fourth, existing approaches inadequately align cross-modal features because they simply average features while disregarding temporal coupling relation. Motivated by these observations in FSAR, we propose Task-Adapter++, a parameter-efficient dual adaptation framework for both image and text encoder. This method comprehensively incorporates task-specific information for modeling video features and preserves the ordering information in enhanced action descriptions.
@@
To address the drawback of full fine-tuning, we turn to the recently emerged PEFT approach. Originating from the field of NLP [13, 12, 16], PEFT aims to adapt large pre-trained models to downstream applications without fine-tuning all the parameters of the model, while attaining performance comparable to that of full fine-tuning. In the field of computer vision, a series of adapter-tuning works [54, 17] have demonstrated the effectiveness of PEFT for video understanding. Inspired by these works, we argue such methods are particularly well-suited for FSAR, since they enable better trade-offs between leveraging the generalizable prior knowledge of pre-trained models and incorporating domain-specific knowledge from downstream tasks.
On the basis of the existing PEFT, we hope to introduce task-specific adaptation into feature extraction. Therefore, we propose a novel adapter-based framework with task-specific adaptation capabilities. Unlike existing methods that apply task adaptation to extracted features, we uniquely integrate it into the feature extractor itself. As shown in Fig. 1 (b), by reusing the frozen self-attention blocks to perform cross-video task-specific attention, our approach dynamically highlights discriminative task-relevant patterns within the support set. This innovation bridges a critical gap in few-shot learning: balancing the preservation of generalizable pre-trained knowledge with the integration of domain-specific cues, enabling more robust and sample-efficient video understanding.
@@
Thirdly, motivated by the observation that real-world events inherently follow stage-specific sub-actions, we leverage LLM to generate and structurally link three temporally ordered sub-actions (derived from input labels) that describe an action’s beginning, process and end. Unlike prior text-based enhancements that treat descriptions as unordered, our approach enforces sequential dependencies between sub-actions, aligning semantic hierarchies with the temporal structure of videos. Further, we propose semantic order adaptation, a novel method that explicitly models the hierarchical progression of actions through language, displayed in Fig. 1 (c). This enables the semantic branch to complement visual-temporal features with fine-grained, order-aware context—a critical advancement for bridging the modality gap in FSAR.
At last, we propose a fine-grained cross-modal alignment strategy tailored for temporally video features. Unlike prior methods that force global average feature, our method first finely models the adjacent frame and then matches cross-modal features by stages. To be specific, we enhance the similarity calculation that will be performed separately on the three semantic and visual features, shown in Fig 1 (c). This innovation addresses a key gap in FSAR: matching multimodal cues finely within structured temporal stages rather than treating actions as a whole unit.
In conclusion, overall of framework is presented in Section 3 and our contributions can be summarized into the following four points:
@@
  • We extend the semantic information with dynamic sub-action description using LLM and further model the sequential relationship of actions by semantic order adaptation.
  • We propose a fine-grained cross-modal alignment method by modeling visual features, which helps mapping visual features to reside in the same temporal stage as semantic description.
  • We clearly improve the performance over the previous version, with an overall improvement up to 3.4% on 5 benchmarks. And additional experiments fully demonstrate the advanced nature and effectiveness of each component of our Task-adapter++.
In this paper, we extend our previously published conference paper in MM’24 [3] in the following main aspects. (i) We propose a dual adaptation method for few-shot action recognition, which fine-tune the adapters concurrently in both image and text encoders. (ii) Benefiting from the LLM, we extend the labels to more comprehensive and fine-grained sub-action descriptions. This step greatly enriches semantic information on the temporal level. (iii) We further propose an adaptation strategy for the semantic branch, namely semantic order adapter, which better models sequential relationships of sub-actions. (iv) We introduce a fine-grained cross-modal alignment strategy ensuring visual features resides in the same temporal stage with semantic features. (v) We demonstrate each component’s effectiveness by comprehensive ablation experimental analysis and the mechanism of module action is further elucidated. (vi) We improve the SOTA performance over 5 benchmarks.
Figure 3: Illustration of Task-Adapter++. Note that we separately add task adapters and semantic order adapters into the two branches, i.e., image encoder and text encoder. After feature extraction, support and query prototypes are passed to the metric module to compute the classification scores and the query prototypes are sent into a alignment module before the metric computation with semantic features. The final score can be calculated on these two.
@@ 2.1 Few-shot Learning
Few-shot learning aims to learn a model that can quickly adapt to unseen classes using only a few labeled data. Existing works can be divided into three categories: augmentation-based [25, 46], metric-based [55, 2, 33, 39, 50, 7, 21, 62] and optimization-based methods [8, 53]. Augmentation-based methods use auxiliary data to enhance the feature representations or generate extra samples [25] to increase the diversity of the support set. Metric-based methods focus on learning a generalizable feature extractor and classify the query instance by measuring the feature distance. Optimization-based methods aim to learn a good network initialization which can fast adapt to unseen classes with minimal optimization steps. Among these approaches, metric-based methods are commonly adopted thanks to its simplicity and superior performance. Additionally, task-specific learning has been explored in image-domain metric-based works [32, 1, 48, 24, 18] to achieve better few-shot learning performance. FEAT [55] introduces set-to-set function to obtain task-specific visual features for the few-shot learning tasks. Ta-Adapter [59] introduces a novel approach that enhances few-shot image classification by incorporating task-aware CLIP encoders and a collaborative prompt learning strategy, demonstrating superior performance. However, these methods all introduce task-specific modules at the level of well-extracted image features, which is insufficient for more complex video features due to the additional temporal dimension. In this work, we follow the metric-based paradigm and perform a novel adaptation during the feature extraction process to achieve more comprehensive adaptation to video data.
@@ 2.2 Few-shot Action Recognition
In few-shot action recognition, most of the previous works [26, 58, 30, 66, 23, 68, 14, 35, 52, 27, 20] fall into the metric-based methods. To solve the problem of temporal misalignment when comparing different videos, a wide series of temporal alignment modules are proposed. OTAM [4] improves the dynamic time warping algorithm to preserve the temporal ordering inside the video feature similarities. TRX [30] proposes to take the ordered tuples of frames as the least matching unit instead of single frames. CMOT [23] uses the optimal transport algorithm to find the best matching between two video features. HyRSM [43] treats the sequence matching problem as a set matching problem for better alignment. Except for the aforementioned works, there are also studies focusing on spatial relations [60] and multi-modal fusion [9, 40, 47]. However, there is still a lack of research on task-specific learning in few-shot action recognition. In this work, we emphasize the vital importance of task-specific adaptation in identifying the most discriminative spatial-temporal features within the provided video samples.
@@ 2.3 Parameter-Efficient Fine-Tuning
Parameter-Efficient Fine-Tuning (PEFT) is derived from NLP [13, 12, 16], aiming to fine-tune large pre-trained models on downstream tasks in an efficient way while achieving comparable performance to full fine-tuning. CLIP-Adapter [10] adds a learnable bottleneck layer behind the frozen CLIP backbone to learn new features for downstream tasks. AIM [54] reuses the frozen pre-trained attention layer and applies it to the temporal dimension of the input video. SAFE [67] applies PEFT to semantic awareness that enhance the comprehension ability of image classification. In this work, we propose a novel PEFT method to acquire the most discriminative information and order information on two branches of CLIP.
@@ 3.1 Problem definition
Few-shot action recognition aims to learn knowledge from base classes and classify human action on new classes with limited data. To achieve this goal, we adopt an episode-based training manner, which can be performed as $C$-way $K$-shot. In each episode, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from the training set $\mathcal{D}_{tr}$ for each episode during training. All the $C\times K$ video samples form the support set $\mathcal{S_{tr}}=\{x_{i}\}_{i=1}^{C\times K}$. Additionally, there is a query set $\mathcal{Q}_{tr}=\{x_{i}\}_{i=1}^{Q_{tr}}$ from the same $C$ action classes. The learning objective of each task is to classify each query video $q\in\mathcal{Q}_{tr}$ into one of the $C$ classes. In test stage, we sample $C$-way $K$-shot task from testing set $\mathcal{D}_{ts}$ and classify the query video $q\in\mathcal{Q}_{ts}$ into one of the $C$ classes with frozen parameters.
Figure 4: In task-specific adaptation, we introduce task adapter after T-MSA and S-MSA to enhance the task-specific information for the few-shot action recognition. This part illustrates the computational process of the image encoder given a 2-way 1-shot learning task with two query videos in the query set.
@@ 3.2 Framework
The architecture of whole Task-adapter++ is shown in Fig. 3, which reveal the main framework of our dual adaptation. The visual branch is shown in Fig. 4 where we introduce task specific adaptation to pay attention to the most discriminative video information. And Fig. 5 shows the semantic branch which model the sequential information between sub-actions by our proposed semantic order adaptation.
@@ 3.3 Adapting ViT to Action Recognition
Given an RGB video sample $v\in\mathbb{R}^{T\times H\times W\times 3}$, a vanilla ViT takes the temporal dimension $T$ as the batch dimension $B$, i.e., there is no exchange of information between frames. Specifically, ViT divides each frame into $N=HW/P^{2}$ patches, where $P$ is the patch size, and transforms the patches into $D$ dimensional patch embeddings $v_{p}\in\mathbb{R}^{T\times N\times D}$. A learnable [class] token is prepended to the patch embeddings $v_{p}$ as $v_{0}\in\mathbb{R}^{T\times(N+1)\times D}$. To retain position information, a learnable positional embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ is added to $v_{0}$ as $z_{0}=v_{0}+E_{pos}$, where the $E_{pos}$ is broadcast to $T$ dimension to be shared among different frames and $z_{0}$ is the final input of the stacked ViT blocks. Each ViT block consists of LayerNorm (LN), multi-head self-attention (MSA) and MLP layers, with residual connections after each block, as shown in Fig. 2 (a). The computation process during each ViT block can be given by:
@@
Inspired by PEFT techniques, AIM tries to adapt pre-trained image model to video action recognition with minimal fine-tuning by freezing the original model and only fine-tuning additionally introduced adapters shown in Fig. 2 (b). Moreover, AIM proposes to reuse the pre-trained MSA layer to do temporal modeling which is deficient in the vanilla ViT. As shown in Fig. 2 (c), a shared frozen temporal MSA (T-MSA) layer is prepended to the original spatial MSA (S-MSA) sequentially to exchange the temporal information between frames for temporal modeling, which can be achieved easily by permuting the patch dimension and the temporal dimension. The tunable Adapter is added after each MSA layer and parallel to the MLP layer to achieve temporal, spatial and joint adaptation as follows:
@@
Figure 5: In semantic order adaptation, we introduce semantic order adapter after MSA to model the sequential information of sub-actions. This part illustrates the computational process of the text encoder given a 2-way 1-shot learning task.

### v2

@@ Abstract
Few-shot action recognition, Parameter-efficient fine-tuning, Task-specific adaptation, Semantic order adaptation
journal: Pattern Recognition
@@
[…]
{graphicalabstract}
{highlights}
We propose Task-Adapter++, a parameter-efficient dual adaptation framework for few-shot action recognition, which alleviates catastrophic forgetting and overfitting induced by full fine-tuning.
For the visual branch, we propose Task-Adapter to extract discriminative task-specific information during feature extraction.
For the semantic branch, we propose semantic order adaptation, which utilizes dynamic descriptions to model the sequential relationships among sub-actions, thereby enhancing the performance of stage-wise cross-modal matching.
Task-Adapter++ achieves state-of-the-art performance across five commonly-used few-shot action recognition benchmarks.
Figure 1: Comparison of different paradigms. (a) Existing methods fully fine-tune the feature extractor and combine it with temporal alignment module. (b) Existing methods extract semantic features from static enhanced descriptions and calculate the score by global matching. (c) For the image encoder, we freeze the backbone and apply task-specific adaptation to extract the most discriminative information. (d) For the text encoder, we use semantic order adapter to model sequential information within dynamic sub-actions and further calculate the score by stage with aligned visual features.
@@ 1 Introduction
With the advent of large-scale pre-trained models [1, 2], researchers are now able to exploit the knowledge these models have acquired during pre-training and apply it to downstream tasks. Accordingly, there is an increasing focus on fine-tuning large pre-trained models for specific tasks [3, 4, 5, 6]. In video understanding, existing works typically fine-tune pre-trained models on well-labeled video datasets [7, 8], where the training and testing categories are identical. However, in realistic scenarios, collecting sufficient and well-labeled video samples is inherently challenging. As a result, few-shot learning has garnered increasing attention [9, 10, 11, 12, 13].
A common practice in few-shot action recognition is to adopt a metric-based meta-learning strategy, where classification probabilities are computed based on the distance between query video features and support prototypes. Prior works [14, 15, 16, 17] achieve this by fine-tuning large image models, such as CLIP [1], on video datasets alongside independent temporal modules, as shown in Fig. 1 (a). Beyond visual-centric approaches, recent studies have explored multimodal frameworks for few-shot learning by integrating semantic information. For instance, works like [18, 19] utilize the zero-shot classification ability of CLIP by video-text matching, significantly improving the accuracy of few-shot classification. Some works [20, 21, 22] additionally employ large language models (LLMs) or vision-language models (VLMs) to expand label information, enriching the attribute information of actions or objects, as shown in Fig. 1 (b).
@@
Although previous studies have made progress, several limitations remain. First, full fine-tuning erodes the model’s inherent prior knowledge, thereby degrading its ability to generalize to unseen tasks. Second, the importance of incorporating task-specific adaptation into spatio-temporal feature extraction is often overlooked, which leads to the neglect of the most discriminative information in each task. Third, the enhanced label information is typically treated as a static description, ignoring the different stages of action occurrence. Due to the absence of fine-grained temporal descriptions and semantic order modeling, it is difficult to fully exploit CLIP’s zero-shot classification capabilities in video understanding. Fourth, existing approaches inadequately align cross-modal features by global matching, which disregards the temporal coupling between visual and semantic features. Motivated by these observations, we introduce Task-Adapter++, a framework that optimally caters to few-shot action recognition (FSAR) through parameter-efficient dual adaptation. Our method comprehensively incorporates task-specific information during video feature extraction, and preserves semantic order information throughout both semantic modeling and cross-modal alignment.
@@
Specifically, to address the first drawback of full fine-tuning, we turn to the recently emerged parameter-efficient finetuning (PEFT) approach. Originating from the field of NLP, PEFT [23, 24, 25] aims to adapt large pre-trained models to downstream applications without fine-tuning all the parameters of the model, while attaining performance comparable to that of full fine-tuning. Inspired by these works, we argue such methods, which only fine-tune a small number of additional adapter parameters, are particularly well-suited for few-shot action recognition, since they balance the preservation of generalizable pre-trained knowledge with the integration of domain-specific cues, enabling more robust and data-efficient video understanding.
As for the second drawback, existing adapter methods in video domain mainly focus on temporal modeling and lack the consideration of enhancing task-specific features for the few-shot learning tasks. Regarding human cognition, a common consensus is that the most discriminative features vary for different target tasks. For example, in distinguishing between “High Jump” and “Long Jump”, the body movements performed by the actor is more crucial, whereas in distinguishing between “High Jump” and “Pole Vault”, it is more significant whether the actor is holding a long pole. Extracting task-specific features has been a key issue in metric based few-shot image classification  [12, 26], while neglected in FSAR. Although a few works [16, 27] consider task-specific adaptation, they solely apply task-specific methods at the level of well-extracted features, which hinders the discovery of the most discriminative information for a given few-shot learning task. In this work, we propose a novel adapter-based method that brings the capability of task-specific adaptation directly into the process of video feature extraction. As shown in Fig. 1 (c), by reusing the frozen self-attention blocks with learnable adapters to perform cross-video task-specific attention, our approach dynamically highlights the most discriminative task-relevant patterns within each given task.
@@
Thirdly, motivated by the observation that real-world events inherently follow stage-specific sub-actions, we leverage LLMs to generate and structurally link temporally ordered sub-actions that describe an action’s beginning, process and end. Unlike prior text-based enhancements that regard descriptions as static constructs, our approach enforces sequential dependencies between sub-actions, aligning semantic sub-action with the temporal structure of videos. Furthermore, we propose semantic order adaptation, a novel method that explicitly models the sub-action processes underlying complex actions through language, as displayed in Fig. 1 (d). This enables the semantic branch to provide fine-grained, order-aware context for the subsequent matching.
Finally, when matching visual and semantic features, to overcome the limitations of global matching and fully take advantage of the temporal coupling between visual and semantic features, we propose a novel fine-grained cross-modal alignment strategy. This strategy aligns video features with semantic features across different stages. Specifically, after fine-grained temporal modeling, which is realized by an attention mechanism, we segment the visual features to directly correspond with the semantic features representing the beginning, process, and end stages. As shown in Fig. 1 (d), the final cross-modal matching score is derived from this fine-grained, segment-wise alignment across different temporal stages.
Figure 2: Illustration of Task-Adapter++. We first augment the labels with sub-action descriptions using LLMs, and then independently integrate task adapters and semantic order adapters into the visual and semantic branches. After feature extraction, the support and query features are input into a metric module to compute classification scores, while the query features are further processed through a cross-modal alignment module prior to cross-modal video-text matching. The final score is obtained by combining the outputs from video matching and video-text matching.
As a whole, our framework consists of two branches and computes the matching score accordingly: a visual branch with an image encoder enhanced with task-specific adaptations, and a semantic branch with a text encoder incorporating semantic order adaptations. As shown in Fig. 2, the visual matching score is computed based on the query video features and the support video prototypes extracted by the visual branch, while the cross-modal matching score is derived from the semantic features extracted by the semantic branch and the query video features after cross-modal alignment. The final score is obtained by multiplying these two parts.
Our contributions can be summarized as follows:
  • 1.
@@
  • 3.
    We extend the semantic information with dynamic sub-action descriptions using LLMs and further model the sequential relationship of actions by semantic order adaptation. After that, we propose a fine-grained cross-modal alignment method which helps matching visual features with semantic features by stages.
  • 4.
    Extensive experiments fully demonstrate the superiority of our Task-Adapter++, establishing new state-of-the-art results on five standard few-shot action recognition benchmarks, and validating the effectiveness of each proposed component.
In this paper, we extend our previously published conference paper in MM’24 [28] in the following aspects. (i) We propose a novel dual adaptation method for few-shot action recognition, which fine-tunes the adapters concurrently in both image and text encoders. (ii) Benefiting from the LLMs, we extend the labels to more comprehensive and fine-grained sub-action descriptions. This significantly enriches the semantic information at the temporal level. (iii) We further propose a novel adaptation strategy for the semantic branch, namely semantic order adapter, which better models the sequential relationships of sub-actions. (iv) We introduce a fine-grained cross-modal alignment strategy ensuring visual features reside in the same temporal stage as semantic features for better cross-modal feature matching. (v) We conduct comprehensive ablation studies to demonstrate the effectiveness of each proposed component of our Task-Adapter++. The underlying mechanism of each module is further elucidated, providing deeper insights into their contributions. (vi) We significantly boost the performance over the previous version, with an improvement up to 3.4% on 5 benchmarks.
@@ 2.1 Few-shot Learning
Few-shot learning aims to learn a model that can quickly adapt to unseen classes using only a few labeled data. Existing works can be divided into three categories: augmentation-based [29, 30], metric-based [12, 13, 31, 32, 33, 34, 35, 36] and optimization-based methods [37, 38]. Augmentation-based methods use auxiliary data to enhance the feature representations or generate extra samples [29] to increase the diversity of the support set. Metric-based methods focus on learning a generalizable feature extractor and classify the query instance by measuring the feature distance. Optimization-based methods aim to learn a good network initialization which can fast adapt to unseen classes with minimal optimization steps. Among these approaches, metric-based methods are commonly adopted thanks to its simplicity and superior performance. Additionally, task-specific learning has been explored in image-domain metric-based works [26, 39, 40, 41] to achieve better few-shot learning performance. FEAT [12] introduces set-to-set function to obtain task-specific visual features for the few-shot learning tasks. Ta-Adapter [42] introduces a novel approach that enhances few-shot image classification by incorporating task-aware CLIP encoders and a collaborative prompt learning strategy, demonstrating superior performance. However, these methods all introduce task-specific modules at the level of well-extracted image features, which is insufficient for more complex video features due to the additional temporal dimension. In this work, we follow the metric-based paradigm and perform a novel adaptation during the feature extraction process to achieve more comprehensive adaptation to video data.
@@ 2.2 Few-shot Action Recognition
In few-shot action recognition, most of the previous works [15, 17, 43, 44, 45, 46, 47, 48] fall into the metric-based methods. To solve the problem of temporal misalignment when comparing different videos, a wide series of temporal alignment modules are proposed. OTAM [14] improves the dynamic time warping algorithm to preserve the temporal ordering inside the video feature similarities. TRX [15] proposes to take the ordered tuples of frames as the least matching unit instead of single frames. CMOT [45] uses the optimal transport algorithm to find the best matching between two video features. HyRSM [16] treats the sequence matching problem as a set matching problem for better alignment. Except for the aforementioned works, there are also studies focusing on spatial relations [49] and multi-modal fusion [50, 51, 52]. However, there is still a lack of research on task-specific learning in few-shot action recognition. In this work, we emphasize the vital importance of task-specific adaptation in identifying the most discriminative spatio-temporal features within the provided video samples.
@@ 2.3 Parameter-Efficient Fine-Tuning
Parameter-Efficient Fine-Tuning (PEFT) is derived from NLP [23, 24, 25], aiming to fine-tune large pre-trained models on downstream tasks in an efficient way while achieving comparable performance to full fine-tuning. CLIP-Adapter [3] adds a learnable bottleneck layer behind the frozen CLIP backbone to learn new features for downstream tasks. AIM [5] reuses the frozen pre-trained attention layer and applies it to the temporal dimension of the input video. SAFE [53] applies PEFT to semantic awareness that enhance the comprehension ability of image classification. However, in the FSAR task, no work has ever considered extracting task-specific visual features and modeling the sequential information of sub-actions using PEFT. In this work, we propose a novel PEFT method to acquire the most discriminative information and order information on two branches of CLIP for FSAR.
@@ 3.1 Problem definition
Few-shot action recognition aims to learn knowledge from base classes and classify human action on new classes with limited data. To achieve this goal, we adopt an episode-based training manner, which can be performed as $C$-way $K$-shot. The video datasets are partitioned into nonoverlapping training and testing splits, denoted as $\mathcal{D}_{tr}$ and $\mathcal{D}_{ts}$, respectively. In each episode, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from $\mathcal{D}_{tr}$ for each episode during training. All the $C\times K$ video samples form the support set $\mathcal{S}_{tr}=\{x_{i}\}_{i=1}^{C\times K}$. Additionally, there is a query set $\mathcal{Q}_{tr}=\{x_{i}\}_{i=1}^{Q_{tr}}$ from the same $C$ action classes. The learning objective of each task is to classify each query video $q\in\mathcal{Q}_{tr}$ into one of the $C$ classes. In test stage, we sample $C$-way $K$-shot task from $\mathcal{D}_{ts}$ and classify the query video $q\in\mathcal{Q}_{ts}$ into one of the $C$ classes with frozen parameters.
Figure 3: Illustration of Task-Adapter in visual branch. For task-specific adaptation, we introduce task adapter after T-MSA and S-MSA to enhance the task-specific information for the few-shot action recognition. This part illustrates the computational process of the image encoder given a 2-way 1-shot learning task with two query videos in the query set.
@@ 3.2 Framework
The architecture of the whole Task-Adapter++ is shown in Fig. 2, which demonstrate our dual adaptation method. The visual branch is shown in Fig. 3 where we introduce task-specific adaptation to pay attention to the most discriminative video information in a given task. And Fig. 5 shows the semantic branch which model the sequential information between sub-actions by our proposed semantic order adaptation. The classification score consists of two parts: a metric score derived from the support and query features, and a cross-modal matching score computed from the query and semantic features. The details of each component are described as follows.
@@ 3.3 Adapting ViT to Action Recognition
Given an RGB video sample $v\in\mathbb{R}^{T\times H\times W\times 3}$, a vanilla ViT takes the temporal dimension $T$ as the batch dimension $B$, i.e., there is no exchange of information between frames. Specifically, ViT divides each frame into $N=HW/P^{2}$ patches, where $P$ is the patch size, and transforms the patches into $D$ dimensional patch embeddings $v_{p}\in\mathbb{R}^{T\times N\times D}$. A learnable [class] token is prepended to the patch embeddings $v_{p}$ as $v_{0}\in\mathbb{R}^{T\times(N+1)\times D}$. To retain position information, a learnable positional embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ is added to $v_{0}$ as $z_{0}=v_{0}+E_{pos}$, where the $E_{pos}$ is broadcast to $T$ dimension to be shared among different frames and $z_{0}$ is the final input of the stacked ViT blocks. Each ViT block consists of LayerNorm (LN), multi-head self-attention (MSA) and MLP layers, with residual connections after each block, as shown in Fig. 4 (b). The computation process during each ViT block can be given by:
@@
Inspired by PEFT techniques, AIM tries to adapt pre-trained image model to video action recognition with minimal fine-tuning by freezing the original model and only fine-tuning additionally introduced adapters shown in Fig. 4 (c). Moreover, AIM proposes to reuse the pre-trained MSA layer to do temporal modeling which is deficient in the vanilla ViT. As shown in Fig. 4 (a), a shared frozen temporal MSA (T-MSA) layer is prepended to the original spatial MSA (S-MSA) sequentially to exchange the temporal information between frames for temporal modeling, which can be achieved easily by permuting the patch dimension and the temporal dimension. The tunable adapter is added after each MSA layer and parallel to the MLP layer to achieve temporal, spatial and joint adaptation as follows: